Find the finite population correction factor given the following:


1. N=200, n=10

2. N=2000, n=10

3. N=400, n=40

4. N=500, n=10

5. N=200, n=20


1
Expert's answer
2021-03-31T14:09:08-0400

"FPC=\sqrt{\frac{N-n}{N-1}}."

  1. "FPC=\sqrt{\frac{200-10}{200-1}}=0.9771."
  2. "FPC=\sqrt{\frac{2000-10}{2000-1}}=0.9977."
  3. "FPC=\sqrt{\frac{400-40}{400-1}}=0.8494."
  4. "FPC=\sqrt{\frac{500-10}{500-1}}=0.9909."
  5. "FPC=\sqrt{\frac{200-20}{200-1}}=0.9511."